OpenStudy (anonymous):

Solve each system. @danJS x^2+y^2+2x+ 2y = 0 x^2+y^2+4x+6y+12 = 0

OpenStudy (phebe):

ill answer them for u

OpenStudy (anonymous):

ok, thanks

OpenStudy (phebe):

for number 1 Simplify x2+2x+y2 + 2y

OpenStudy (phebe):

welcome

OpenStudy (phebe):

for integer solution answer is x = -2, y = -2

OpenStudy (phebe):

Answer for 1 is x^2+2 x+y (y+2) = 0 when simplifyed okay

OpenStudy (phebe):

im doing number 2 now okay

OpenStudy (anonymous):

ok

OpenStudy (phebe):

Integer soulution is x = -3, y = -3

OpenStudy (phebe):

Answer is (x+2)^2+(y+3)^2-1 = 0

OpenStudy (anonymous):

thank you for your help!

OpenStudy (phebe):

welcome ohh medal plzz

OpenStudy (phebe):

need anything else?

OpenStudy (phebe):

well in order to get your answer you basically just have to simplify

OpenStudy (phebe):

an its easier from there

OpenStudy (perl):

there are actually 2 intersection points x = -2 and y = -2 x = -6/5, y = -12/5

OpenStudy (anonymous):

So like, who's right?

OpenStudy (phebe):

Both we got the same answer basically